OpenStudy (anonymous):

solve the system of equations by elimination 4x-2y=0 3x+3y=9

OpenStudy (anonymous):

multiply the first equation by 3, the second one by 2, then add the equations... what do you get as a result?

OpenStudy (anonymous):

there are many ways you can do elimination but with what i described, i eliminated the "y" variable....

OpenStudy (anonymous):

12x-y=0 12x+12y=9

OpenStudy (anonymous):

3(4x-2y=0) -4(3x+3y=9) then multiply across and you get 12x-6y=0 -12x-12y=-36 and do the elimination (:

OpenStudy (anonymous):

notice in @BryanaGarcia 's elimination, she eliminated the "x" variable....

OpenStudy (anonymous):

yea. so what would be left is -18y=-36

OpenStudy (anonymous):

yep... then solve for x... in my method, you'll get the x first then you can solve for y...

OpenStudy (anonymous):

y=18